OpenStudy (anonymous):

Amanda received a discount of 30 % off of the price of a portrait product that costs $75. How much money did Amanda save?

OpenStudy (immanuelv):

75*(.3 ) = 22.5 then u do 75-22.5 = 52.5

OpenStudy (immanuelv):

she saved $22.5

OpenStudy (anonymous):

thanx

OpenStudy (immanuelv):

np...ask me fi u have more questions :)

OpenStudy (anonymous):

yes i do lol. many more. If a salesperson places telephone calls to 80 potential customers, speaks with 60% of them, and books appointments with 50% of the customers she speaks with, how many customers does she book?

OpenStudy (immanuelv):

she speaks with 80* (.6) = 48 and books with 80* (.5) = 40

OpenStudy (immanuelv):

srry she books with 48 (.5) = 24

OpenStudy (anonymous):

ok that what i had 24

OpenStudy (immanuelv):

speaks with 48 people and she books with 24 :) srry for my mistake!

OpenStudy (anonymous):

its ok another one..If a special portrait package offer consists of 2 8x10-sized portraits, 2 5x7-sized portraits, and 8 wallet-sizes portraits, what fraction of the portraits are 5x7-sized?

OpenStudy (immanuelv):

hmm....there are 3 types of pics...... and u can get 10 pics...... so to find portraits u just divide 2/10 which is (2/10) or .2

OpenStudy (anonymous):

so is .2

OpenStudy (immanuelv):

yea

OpenStudy (immanuelv):

fraction is 2/10